OpenStudy (anonymous):

let f(x) be -11xsinxcosx, find f'(3pi/2)

OpenStudy (jwt625):

\[f \prime(x)=-11\sin (x)\cos (x)-11x(\cos (x)^2-\sin (x)^2)\]

OpenStudy (jwt625):

f'(3pi/2)=33pi/2